OpenStudy (riley):

Need help please. For the demand equation, express the total revenue R as a function of the price p per item. q = −5p + 500 R(p) = ? So if I just set R=p(-5p+500), I can turn it into a quadratic formula . . . and ummm still working on where to go from there . . . >_<

OpenStudy (riley):

Nevermind I figured it out :) R=pq =p(-5p+500) =-5p^2+500p

OpenStudy (riley):

So then I take P=-b/2a and plug that in to figure out p.

OpenStudy (riley):

And end up with p=50, and to get my revenue equation I just use =5p^2+500p

OpenStudy (riley):

and 50 is the maximum for the total revenue xD I think I did it right . . .

OpenStudy (amistre64):

what does the: q = -5p + 500, represent?

OpenStudy (riley):

the quantity

OpenStudy (amistre64):

q = quantity? at a given p, price?

OpenStudy (riley):

yesh

OpenStudy (riley):

But I think I got it, quadratic function is my friend<3

OpenStudy (amistre64):

and revenue = price, p times q quantity sold

OpenStudy (riley):

Unless I did it wrong O_O . . . I don't think I did . . . it makes sense when I go back and rework my math.

OpenStudy (riley):

yes sir

OpenStudy (amistre64):

the quadratic is good; it lets you determine where the maximum revenue with regards to price is then

OpenStudy (amistre64):

you did good

OpenStudy (amistre64):

R(p) = -5p^2 + 500p p = -500/2(-5) = 500/10 = 50

OpenStudy (riley):

yes thats what I did :D

OpenStudy (riley):

Yay maybe theres hope yet for me, I think I'm understanding this now. Tyvm for checking it for me.

OpenStudy (amistre64):

youcould have also found where it zeros out at and get the middle between them

OpenStudy (amistre64):

youre welcome

OpenStudy (riley):

umm how would I find where it zeroes out? o.o

OpenStudy (riley):

do I set it equal to zero?

OpenStudy (amistre64):

factor it: -5p^2 + 500 factors to: p(-5p + 500) = 0 when p=0 and p=100 the middle point being 50

OpenStudy (riley):

ohhhhhhhh, wow that makes sense now too haha omg

OpenStudy (amistre64):

:) good luck

OpenStudy (riley):

thanks again ^.^<3